OpenStudy (anonymous):

Determine the slope between the points (-3, 0) and (0, 5)

OpenStudy (anonymous):

5-0/0-(-3)=5/3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

5-0, 0-(-3) = 5/3